1. For a car travelling with speed v around a curve of radius r.
a) Determine a formula for the angle at which a road should be banked so that no
friction is required.
b) What is this angle for an expressway off-ramp curve of radius 50 m at a design
speed of 50 km/h?
1
Expert's answer
2020-05-19T10:49:54-0400

a) Let's consider the case when the car travels on a frictionless banked road with speed "v" around a curve of radius "r". There are two forces that act on the car: weight, "mg", and the normal force "N". We can resolve the normal force into the horizontal ("Nsin \theta" ) and vertical ("Ncos \theta") components. The horizontal component of the normal force provides the necessary centripetal force, so that if the car has a right speed there is no friction required.

Let's apply the Newton's Second Law of Motion in projections on "x"- and "y"-axis:


"Nsin \theta = \dfrac{mv^2}{r} (1)""Ncos \theta = mg (2)"

Dividing equation (1) by equation (2) we determine the formula for the angle at which a road should be banked so that no friction is required:


"tan \theta = \dfrac{v^2}{rg},""\theta = tan^{-1}(\dfrac{v^2}{rg})."

b) We can find the angle for an expressway off-ramp curve of radius 50 m at a design speed of 50 km/h from the formula obtained in part (a):


"\theta = tan^{-1}(\dfrac{(50 \dfrac{km}{h} \cdot \dfrac{1000m}{1km} \cdot \dfrac{1h}{3600s})^2}{50m \cdot 9.8 \dfrac{m}{s^2}}) = 21.5^{\circ}."

Answer:

"\theta = 21.5^{\circ}."